Fabri Corporation is considering eliminating a department that has an annual contribution margin of $31,000 and $62,000 in annual fixed costs. Of the fixed costs, $15,500 cannot be avoided. The annual financial advantage (disadvantage) for the company of eliminating this department would be:

Multiple Choice

  • ($31,000)

  • $31,000

  • ($15,500)

  • $15,500
Answer #1

Current net operating income=Contribution margin-Fixed cost

=(31000-62000)=$(31,000)

Unavoidable fixed costs after elimination=($15500)

Hence financial advantage would be =(15500)-(31000)

=$15500